Jim has a part time job selling newspapers and magazine subscriptions. One week he earns $31.10 by selling 5 newspapers and 3 magazine subscriptions. The following week he earns $19.75 by selling 3 newspapers and 2 magazine subscriptions. How much does he earn for each newspaper and each magazine subscription?
Answer by checkley77(12844)

5N+3M=31.10 MULTIPLY BY 2.
3N+2M=19.75 MULTIPLY BY -3 & ADD.
10N+6M=62.20
-9N-6M=-59.25
----------------
N=$2.95 COST OF THE NEWSPAPER.
5*2.95+3M=31.10
[14.75]+3M=31.10 SORRY ABOUT THAT MY KEYBOARD SOMETIMES GETS CONFUSED.
3M=31.10-14.75
3M=16.35
M=16.35/3
M=$5.45 COST OF THE MAGAZINE.